A 10 kg mass is initially moving up a smooth incline a 5 m/s. What distance up the incline will it travel before it stops?


1
Expert's answer
2021-02-23T10:03:38-0500
"0.5mv^2=mgh=mgl\sin{30}\\v^2=gl\\5^2=9.8l\\l=2.55\ m"